IF ONLY FERRIC CHLORIDE IS AVAILABLE IN THE MARKET INSTEAD OF FERRIC SULFATE, CALCULATE THE OPTIMAL MASS YOU WOULD NEED TO PURCHASE EVERY DAY
1
Expert's answer
2020-10-08T13:53:48-0400

Molicular formula of FERRIC CHLORIDE is = FeCl3


Molicular weight of FeCl3

(Atomic mass of Fe + atomic mass of chlorine ×3)

= (56+ 35.5×2)

= 56+ 71

= 127 g/mol




Molicular formula of FERRIC SULFATE is = Fe2(SO4)3


Molicular weight of Fe2(SO4)3

(Atomic mass of Fe + atomic mass of chlorine ×3)

= 56×2+(32×3+12×16

= 400 g/mol


Hence,

If only FERRIC CHLORIDE is available in the market instead of FERRIC SULFATE, I need to purchase 127g FERRIC CHLORIDE for every 400 g of FERRIC SULFATE.